F(x,y)={k(1-x^2) find the value of k



1
Expert's answer
2021-03-16T05:03:40-0400

Since f(x) is a p.d.f. "\int_{\infty}^{\infty} f(x)dx=1"


"\implies \int_0^1 k(1-x^2)dx=1"


"\implies k(x-\dfrac{x^3}{3})_0^1=1"


"\implies k(1-\dfrac{1}{3})=1"


"\implies k=\dfrac{3}{2}"